OpenStudy (anonymous):

Please Help Solve for x: 2|x − 3| + 1 = 7 x = 0, x = 6 x = −1, x = 6 x = 0, x = 9 No solutions

OpenStudy (zzr0ck3r):

first we want to isolate the |stuff| term so subtract both sides by 1, then divide both sides by 2 and what do you have?

OpenStudy (zzr0ck3r):

@Trai7 you with me?

OpenStudy (zzr0ck3r):

I have to go but @Jhannybean will finish, using the property that |a|=b implies a=b and -a=b

OpenStudy (anonymous):

Im here

OpenStudy (zzr0ck3r):

what do you have?

OpenStudy (anonymous):

Im not sure im totally lost

OpenStudy (zzr0ck3r):

subtract 1 from both sides, and what do you have?

OpenStudy (jhannybean):

2|x-3|+1=7 2|x-3| = 6 Divide 2 on both sides |x-3|=3 Open up your absolute value x-3 = 3 or x-3 = -3 x = 3+3 or x = -3 +3

OpenStudy (anonymous):